OpenStudy (exoexoexo):

mike can complete a project in 60 minutes, and if mike and walter both work on the project, they can complete it in 40 minutes. how long will it take walter to complete the project by himself?

OpenStudy (anonymous):

Mike =(1/60) project/min Walter =(w) project/min together (1 + 60 w)/60 project /min [(1+60 w)/60 ] 40 = 1 project 1 + 60 w = 6/4 60 w = .5 w = .0083333 1/w = min/project = 120 minutes